The last time Carolina visited the post office, she spent $4.49 to mail letters and packages. Each letter cost .37 to mail, while each packages cost .88 to mail. If she sent two more letters than packages, how many letters did she mail?
Answer by clearblueskai(22)

The last time Carolina visited the post office, she spent $4.49 to mail letters and packages. Each letter cost .37 to mail, while each packages cost .88 to mail. If she sent two more letters than packages, how many letters did she mail?
let x = number of letters y=number of packages
equations:
.37x+.88y=4.49
x=y+2
.37(y+2)+.88y=4.49
.37y+.74+.88y=4.49
1.25y=3.75
y=3
x=3+2
x=5
no. of letters= 5
no. of packages= 3